A,B,C in order toss a coin. the one to throw a head wins.what are their respective chances of winning assuming that the game may continue indefinetely????????/

Grade:12

1 Answers

AJIT AskiitiansExpert-IITD
13 years ago

Dear  abhishek ,

let the event a THROWS HEAD be denoted by A , not win( THROW A TAIL) by A'. similar for B and C. sp , P(A) =P(A') =1/2 and same for b and c.

P( Awin) = P( A wins in first turn ) + P( A wins in two turns ) +P( A wins in three turns ) +..........  ad.infinity

P(Awin)  = P(A) + P(A'B'C'A) +P( A'B'C'A'B'C'A).....

P(Awin) =  1/2 + (1/2 )^4 + (1/2)^7....

P(Awin) = (1/2)/(1-(1/2)^3)  = 4/7

P(B wins) = P(A'B) + P(A'B'C'A'B) +P(A'B'C'A'B'C'A'B)+........

P(BWINS) = (1/2)^2 + (1/2)^5 +(1/2)^8.....

P(B wins) = 2/7

similarly , P(C wins) = 1/7

hope it helps !!!!

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

Askiitians Expert

Ajit Singh Verma IITD